40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 1 16:38:46 2012, steve, Update, SUS, spare SRMU 03 optic moved IMG_1464.JPGIMG_1465.JPGIMG_1466.JPG
    Reply  Wed Aug 1 17:04:22 2012, Jenne, Update, SUS, spare SRMU 03 optic moved 
Message ID: 7071     Entry time: Wed Aug 1 16:38:46 2012     Reply to this: 7072
Author: steve 
Type: Update 
Category: SUS 
Subject: spare SRMU 03 optic moved 

SOS optics prepared to be hanged are moved from the South Flow Bench to S15 Clean Cabinet.

SRMU 03 (1-25-2010) specification summery E080460-05-D,  older vintage SRM 01 and PRM 02 (need more specification)

There was one  NOT MARKED SOS with two broken magnets on its face. This is labeled ???

 

This was done to prepare clean space for TIP-TILT drive- test set up.

The existing cable from 1X5 can reach only to the south end:  from whitening filter to satelite amp.  This will be good for future test of suspensions.

We need to make new cable from 1 X 1 to the south end = 40m long

Attachment 1: IMG_1464.JPG  2.135 MB  | Hide | Hide all | Show all
IMG_1464.JPG
Attachment 2: IMG_1465.JPG  1.844 MB  | Hide | Hide all | Show all
IMG_1465.JPG
Attachment 3: IMG_1466.JPG  2.101 MB  | Show | Hide all | Show all
ELOG V3.1.3-